# Enhanced strong interaction effect in synthetic spin-orbit coupling with mixed spin symmetry

### v2

@@ IV.3 Aligned interaction ($g>0$ and $g_{\downarrow\uparrow}=0$)
Figure 5 reveals the energy difference $E_{j}-E_{0}$ in the intermediate regime of $\Omega$ for two different interaction strengths $g/\hbar\omega\xi=0.48,2$. It is similar to the anti-aligned interaction case $(g_{\downarrow\uparrow}>0)$ that the critical points $\Omega_{\text{c}}$ vary in different interaction strengths. One different feature is that the avoided crossing between the ground state and the first excited states is seen for $g/\hbar\omega\xi=0.48$ when $\Omega/\omega\simeq 25$ but not for $g/\hbar\omega\xi=2$. This is caused by the interplay between the contact interaction and the coherent coupling that have attempt to shift energy in the opposite directions [24]. By increasing $g$, the excited states are pushed up, and accordingly the avoided crossing is shifted up and shades out eventually. The existence of this avoided crossing affects the pseudo-spin population as see below.
@@ IV.3.1 Pseudo-spin population
Figure 6: (a) Pseudo-spin population $p_{\text{aligned}}$, $p_{\text{unaligned}}$ in the ground state for $g/\hbar\omega\xi=0.48$ and $g/\hbar\omega\xi=2$. (b) Excess of $p_{\text{aligned}}$ over the population at the limit of $\Omega\to\infty$ as function of $\Omega/\omega$ and $g/\hbar\omega\xi$, i.e. $p_{\text{aligned}}-1/4$. Negative values are not shown for clear presentation.
@@ IV.3.2 Pair correlation
We have computed the pair correlations for different $g$ and $\Omega$. First, let us focus on $g/\hbar\omega\xi=1$ and change $\Omega$ (see the middle panels at $g=1$ in Fig. 7). Overall, the pair correlation behaves similar to the anti-aligned interaction case but is more sensitive to $\Omega$. Even for relatively small coupling, the pair correlation starts deviating from Gaussian distribution. For $\Omega/\omega=5$, one bump appears at the centre ($x=y=0$), and for increasing $\Omega$ four bumps come out. At $\Omega/\omega=26$, where the populations $p_{\text{aligned}}$, $p_{\text{unaligned}}$ approach the extreme values as shown in Fig. 5, a dent emerges at the centre. For increasing $\Omega$ more, a slit at $x=y$ caused by the contact interaction appears, and the effect of SOC vanishes.
Now we fix the coherent coupling strength to $\Omega/\omega=26$ and change the contact interaction strength to $g/\hbar\omega\xi=0.48,2$. For $g/\hbar\omega\xi=0.48$, the dent at $x=y=0$ vanishes, and for $g/\hbar\omega\xi=2$ three bumps emerge in $x>y$ and in $x<y$ respectively, which is similar to the anti-aligned interaction case. The aligned interaction and the anti-aligned interaction give similar effects to the structure of the pair correlations even though they affect the energy spectrum and the pseudo-spin population differently.
@@ V Conclusions
A natural question is whether the same behaviour is seen in the more-particle systems. Our results cannot guarantee but imply that this would be the case for anti-aligned interactions. In these cases, the mixed spin symmetry affects the pseudo-spin populations, and more-particle systems also have the same spin symmetry. The mixed spin symmetric states may act in the same way. Furthermore, we have studied only contact interactions. The fact that the anti-symmetric pseudo-spin states do not feel the interaction depends on the kind of the interaction, and the results we have shown may change with different interaction types.
@@ Appendix A Justification of cut off $M$
In Sec. IV, we set the SOC strength $k\xi=4$ and the cut off $M=50$. Here, we show that the cut off $M$ is large enough. As discussed in Sec. III.2, the cut off $M$ is the number of the eigenstates of the harmonic oscillator we take. The SOC couples these eigenstates such that the momentum of up spin or down spin is boosted. Although larger SOC needs larger $M$, the deviation from the results when $k=0$ is negligible if $M$ is large enough.
Without coherent coupling $\Omega$, the Hamiltonian (3) can be diagonalised in the basis of $\sigma_{z}$, and the lowest energy can be obtained anaalytically and is the same as when $k=0$. For instance, the kinetic term of the single particle Hamiltonian (4) can be written as $\sum_{j=1}^{N}(p_{j}-\hbar k\sigma_{z})^{2}/2m$. The SOC acts as a moving frame, and the lowest energy does not change. Thus, by looking at the lowest energy for different $k$, we see whether the cut off $M$ is large enough. Figure 8 shows the energy computed with $M=50$ as function of $k$. For large $k$, the energy obtained numerically deviates from the energy with $k=0$. For strong interactions, the deviation is worse because strong interactions also couple highly excited eigenstates. For $k\xi=4$, the error is small and at worst about $2.7\%$ for $g/\hbar\omega\xi=g/\hbar\omega\xi=10$, and we have taken $k\xi=4$ in the main text.
@@
Appendix B Population of mixed spin symmetry
Figure 9: Example of transforming the second quantization representation to the first quantization representation. The left hand shows a state of three indistinguishable particles, and the right hand shows superposition between three states of three distinguishable particles. These are summed up due to bosons.
In this work, we investigate the population of mixed spin symmetry and here explain how to compute it briefly. The mixed spin symmetry basis are described in the first quantization representation in Eq. (II). On the other hand, we adopt the second quantization representation for exact diagonalization (see Sec. III.2). Therefore, we transform the second quantization representation of the Hamiltonians (7)(III.2) to the first quantization representation, i.e., label the particles and treat the system as distinguishable particles (see Fig. 9). This enables us to obtain the quantum number and the pseudo-spin state of each particles, i.e., have |$\ket{n_{1},\sigma_{1};n_{2},\sigma_{2};n_{3},\sigma_{3}}$ for $n_{1},n_{2},n_{3}=0,1,\ldots,M-1$ and $\sigma_{1},\sigma_{2},\sigma_{3}=\downarrow,\uparrow$. We transform the pseudo-spin basis using $\downarrow$,$\uparrow$ , ↑ to the classified pseudo-spin basis $S_{j}$, $M_{j}$ by using the relations (II)(II). For more concrete details, see our code [19].

### v3

@@ IV.3 Aligned interaction ($g>0$ and $g_{\downarrow\uparrow}=0$)
Figure 6 reveals the energy difference $E_{j}-E_{0}$ in the intermediate regime of $\Omega$ for two different interaction strengths $g/\gamma=0.48,2$. It is similar to the anti-aligned interaction case $(g_{\downarrow\uparrow}>0)$ that the critical points $\Omega_{\text{c}}$ vary in different interaction strengths. One different feature is that the avoided crossing between the ground state and the first excited states is seen for $g/\gamma=0.48$ when $\Omega/\omega\simeq 25$ but not for $g/\gamma=2$. This is caused by the interplay between the contact interaction and the coherent coupling that have attempt to shift energy in the opposite directions [29]. By increasing $g$, the excited states are pushed up, and accordingly the avoided crossing is shifted up and shades out eventually. The existence of this avoided crossing affects the pseudo-spin population as see below.
@@ IV.3.1 Pseudo-spin population
Figure 7: (a) Pseudo-spin population $p_{\text{aligned}}$, $p_{\text{unaligned}}$ in the ground state for $g/\gamma=0.48$ and $g/\gamma=2$. (b) Excess of $p_{\text{aligned}}$ over the population at the limit of $\Omega\to\infty$ as function of $\Omega/\omega$ and $g/\gamma$, i.e. $p_{\text{aligned}}-1/4$. Negative values are not shown for clear presentation.
@@ IV.3.2 Pair correlation
We have computed the pair correlations for different $g$ and $\Omega$. First, let us focus on $g/\gamma=1$ and change $\Omega$ (see the middle panels at $g=1$ in Fig. 8). Overall, the pair correlation behaves similar to the anti-aligned interaction case but is more sensitive to $\Omega$. Even for relatively small coupling, the pair correlation starts deviating from Gaussian distribution. For $\Omega/\omega=5$, one bump appears at the centre ($x=y=0$), and for increasing $\Omega$ four bumps come out. At $\Omega/\omega=26$, where the populations $p_{\text{aligned}}$, $p_{\text{unaligned}}$ approach the extreme values as shown in Fig. 6, a dent emerges at the centre. For increasing $\Omega$ more, a slit at $x=y$ caused by the contact interaction appears, and the effect of SOC vanishes.
Now we fix the coherent coupling strength to $\Omega/\omega=26$ and change the contact interaction strength to $g/\gamma=0.48,2$. For $g/\gamma=0.48$, the dent at $x=y=0$ vanishes, and for $g/\gamma=2$ three bumps emerge in $x>y$ and in $x<y$ respectively, which is similar to the anti-aligned interaction case. The aligned interaction and the anti-aligned interaction give similar effects to the structure of the pair correlations even though they affect the energy spectrum and the pseudo-spin population differently.
@@ V Conclusions
A natural question is whether the same behaviour is seen in the more-particle systems. Our results cannot guarantee but imply that this would be the case for anti-aligned interactions. In these cases, the mixed spin symmetry affects the pseudo-spin populations, and more-particle systems also have the same spin symmetry. The mixed spin symmetric states may act in the same way. Furthermore, the results we have shown may change by inducing different types of interactions. For example, here we have studied only repulsive contact interactions between two particles where anti-symmetric pseudo-spin states do not feel the interaction. This property would be different if we were also to consider three-body interactions [30].
Additionally, we have focussed on one dimension, but it is possible to induce SOC in two dimensions [20, 31]. There are some differences in SOC in one dimension and two dimensions. In two dimensions, there are the Rashba type and the Dresselhaus type of SOC [32], and they are given by $\left(\hat{p}_{x}\hat{\sigma}_{x}\pm\hat{p}_{y}\hat{\sigma}_{y}\right)\hbar k/m$, respectively. By inducing both types with equal weight, the SOC term is given by $2\hat{p}_{x}\hat{\sigma}_{x}\hbar k/m$, which is essentially the same term considered in one dimension after rotating the frame. Therefore, the results we have shown would be observed in two dimensions when such type of SOC is applied [3]. Otherwise, unique effects to two dimensions could be seen [20, 32]. Also, we have compared our system with a TG gas in finding qualitative similarity of the SOC system to the strongly interacting system. TG gas is unique to one dimension and does not exist in higher dimensions. It is an interesting question whether such enhancement of strong interaction effects is still observed in higher dimensions.
@@ Appendix A Justification of cut off $M$
In Sec. IV, we set the SOC strength $k\ell=4$ and the cut off $M=50$. Here, we show that the cut off $M$ is large enough. As discussed in Sec. III.2, the cut off $M$ is the number of the eigenstates of the harmonic oscillator we take. The SOC couples these eigenstates such that the momentum of up spin or down spin is boosted. Although larger SOC needs larger $M$, the deviation from the results when $k=0$ is negligible if $M$ is large enough.
Without coherent coupling $\Omega$, the Hamiltonian (3) can be diagonalised in the basis of $\sigma_{z}$, and the lowest energy can be obtained anaalytically and is the same as when $k=0$. For instance, the kinetic term of the single particle Hamiltonian (4) can be written as $\sum_{j=1}^{N}(p_{j}-\hbar k\sigma_{z})^{2}/2m$. The SOC acts as a moving frame, and the lowest energy does not change. Thus, by looking at the lowest energy for different $k$, we see whether the cut off $M$ is large enough. Figure 9 shows the energy computed with $M=50$ as function of $k$. For large $k$, the energy obtained numerically deviates from the energy with $k=0$. For strong interactions, the deviation is worse because strong interactions also couple highly excited eigenstates. For $k\ell=4$, the error is small and at worst about $2.7\%$ for $g/\gamma=g_{\downarrow\uparrow}/\gamma=10$, and we have taken $k\ell=4$ in the main text.